Find the equation to the straight line which passes through the point (4,-5) and which is parallel to the line 3x+4y+5=0

Grade:10

1 Answers

Aditya Gupta
4 years ago
by slope point form of line, we can write 
y – ( – 5)= m(x – 4), but as this line is parallel to 3x+4y+5=0, so its slope m would equal to the slope of 3x+4y+5=0. clearly, slope of 3x+4y+5=0 is – ¾.
so eqn becomes
y+5= – (3/4)*(x – 4)
simplify
3x+4y+8= 0
kindly approve yar samaira :=))